A medical researcher is studying the effects of a drug on blood pressure. Subjects in the study have their blood pressure taken at the beginning of the study. After being on the medication for 4 weeks, their blood pressure is taken again. The change in blood pressure is recorded and used in doing the hypothesis test. Change: Final Blood Pressure - Initial Blood Pressure The researcher wants to know if there is evidence that the drug reduces blood pressure. At the end of 4 weeks, 36 subjects in the study had an average change in blood pressure of -2.0 with a standard deviation of 4.5. Find the p -value for the hypothesis test. Your answer should be rounded to 4 decimal places.
Answer #1

Ho :   µd=   0
Ha :   µd <   0

mean of difference ,    D̅ =    -2.000                  
                          
std dev of difference , Sd =        4.5000                  
                          
std error , SE = Sd / √n =    4.5000   / √   36   =   0.7500      
                          
t-statistic = (D̅ - µd)/SE = (   -2   -   0   ) /    0.7500   =   -2.6667

Degree of freedom, DF=   n - 1 =    35  
  
p-value =        0.0058   [excel function: =t.dist(t-stat,df) ]